40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 27 21:48:31 2014, Jenne, Update, LSC, X green broadband PD not working???!? ALS_outofloop_27May2013_2.pdf
    Reply  Wed May 28 17:51:48 2014, manasa, Update, LSC, X green broadband PD NOT working 
       Reply  Wed May 28 19:15:38 2014, Koji, Update, LSC, X green broadband PD NOT working 
Message ID: 10001     Entry time: Wed May 28 19:15:38 2014     In reply to: 10000
Author: Koji 
Type: Update 
Category: LSC 
Subject: X green broadband PD NOT working 

If the PD is the suspect, just pull it from the table and bring it to the PD testing setup.

The transimpedance of the PD should be ~2000 Ohm for both of the RF and DC outputs.

The test setup gives you the systematic opportunity for examination of the signal line.
Check the signal level with the active probe.

Once the broken component is found replace it. You are supposed to have the replacement
components on the blue tower.

ELOG V3.1.3-